OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A. {–12, 5} B. {–11.1, –2.9} C. {–5.6, –1.4} D. {–3.9, 12.1}

OpenStudy (campbell_st):

do you know the general quadratic formula...?

OpenStudy (campbell_st):

ok.... so you don't know the formula.... or you just want answers to your homework quiz...?

OpenStudy (xmoses1):

If you know the formula then the question is relatively easy ^^

OpenStudy (anonymous):

sorry i didnt read that i have the math book in front of me but im not understanding it

OpenStudy (campbell_st):

ok... so the book would say something like... for a quadratic equation \[ax^2 + bx + c = 0\] the general quadratic formula is \[x=\frac{-b \pm \sqrt{b^2 - 4 \times a \times c}}{2 \times a}\] in your question a = 1, b = 7 and c = 8 just substitute them.... then calculate the 2 answers...

OpenStudy (campbell_st):

well when you substitute you get \[x =\frac{ -7 \pm \sqrt{49 - 32}}{2} = \frac{ -7 \pm \sqrt{17}}{2}\] so the 2 solutions are \[x = \frac{-7 + \sqrt{17}}{2}.....and.......x = \frac{-7-\sqrt{17}}{2}\] that may help you find the solutions...

OpenStudy (anonymous):

i actually did all that but i divided them into eachoter 17 and 2 thats how i got 8.5

OpenStudy (campbell_st):

well there is an answer is the list that's really close to my mental arithmetic solutions